OpenStudy (amberofficial1):

a ball is dropped from a cliff and reaches a speed of 29.4 m/s after 3.0 seconds what is the acceleration?

OpenStudy (mrnood):

the equation you need is v= u + at v is final velocity (=29.4 m/s) u is initial velocity (=0 because the ball is 'dropped') t = 3s so you can put those numbers into the equation and work out a